# Duals and time-reversed flows of generalized Ornstein-Uhlenbeck processes

### v3

@@ 4 The time-reversed flow of the generalized Ornstein-Uhlenbeck process
Let $(V_{s}^{x})_{s\geq 0}$, $x\in\mathbb{R}$, be the solution of the SDE (2.1) for the bivariate Lévy process $(U_{s},L_{s})_{s\geq 0}$ with $\Delta U_{s}>-1$ for all $s\geq 0$. Fix $t>0$ and consider the stochastic process $(R_{s}^{y})_{s\in[0,t]}$, $y\in\mathbb{R}$, corresponding to the reversed stochastic flow defined in (4.4). Then $(R_{s}^{y})_{s\in[0,t]}$, $y\in\mathbb{R}$, is Siegmund-dual to $(V_{s}^{x})_{s\in[0,t]}$, $x\in\mathbb{R}$.
@@
The processes $(T,\eta)$ appearing in Theorem 4.1 and $(W,K)$ appearing in Theorem 3.5 are given for $s\in[0,t]$ by (compare also (4.9) and (2.4))
@@
$K_{s}$ $=-\eta_{s}=-L_{s}+\sigma_{U,L}s+\sum_{0<r\leq s}\frac{\Delta U_{r}\Delta L_{r}}{1+\Delta U_{r}}.$
But since the time-reversed Lévy process $(\tilde{U}_{s},\tilde{L}_{s})_{s\in[0,t]}$ is equal in law to $(-U_{s},-L_{s})_{s\in[0,t]}$ we conclude that $(T_{s},\widetilde{\eta}_{s})_{s\in[0,t]}$ and $(W_{s},K_{s})_{s\in[0,t]}$ are equal in law. The claim then follows from Theorems 3.5 and 4.1. ∎
@@
Corollary 4.2 is not surprising, since reversibility of the stochastic flow is intrinsically related to duality, as discussed in various articles, see e.g. [21]. Also in our case, it is possible to deduce Theorem 3.5 directly from Theorem 4.1. To see this, let $(U,L)$, $V^{x}_{s}$ and $R^{y}_{s}$ be as in Corollary 4.2. For $u\in[0,t]$ denote by $\varphi_{u,t}:\mathbb{R}\to\mathbb{R}$ the stochastic flow which transports $x$ when $V_{u}(\omega)=x$ to $V_{t}(\omega)$, i.e. $\varphi_{u,t}(x)=V_{t}(x)$ conditional on $V_{u}=x$ (we suppress the superscript $x$ here and prefer to work with $V_{t}|V_{0}=x$ rather than $V_{t}^{x}$ for the moment). By time-homogeneity of $V$, we have for $s\in[0,t]$
@@
showing that $(R_{s}^{y})_{s\in[0,t]}$, $y\in\mathbb{R}$, is dual to $(V_{s}^{x})_{s\in[0,t]}$, $x\in\mathbb{R}$. By the proof of Corollary 4.2, $(T_{s},\widetilde{\eta}_{s})_{s\in[0,t]}$ is equal in law to $(W_{s},K_{s})_{s\in[0,t]}$, showing that the process given in Theorem 3.5 (call it $\overline{R}_{s}^{y}$ for the moment) is also dual to $R$. Hence we have given another proof of the duality of $(\overline{R}_{s}^{y})$ to $(V_{s}^{x})$ in Theorem 3.5. This last proof is in line with the reasoning given in Sigman and Ryan [21, Cor. 3.1 (1)], who consider $[0,\infty)$-valued time-homogeneous Markov processes with certain properties.

### v4

@@ 4 The inverse flow of the generalized Ornstein-Uhlenbeck process
Let $(V_{s}^{x})_{s\geq 0}$, $x\in\mathbb{R}$, be the solution of the SDE (1.2) for the bivariate Lévy process $(U_{s},L_{s})_{s\geq 0}$ with $\Delta U_{s}>-1$ for all $s\geq 0$. Fix $t>0$ and consider the stochastic process $(R_{s}^{y})_{s\in[0,t]}$, $y\in\mathbb{R}$, corresponding to the inverse stochastic flow defined in (4.4). Then $(R_{s}^{y})_{s\in[0,t]}$, $y\in\mathbb{R}$, is Siegmund dual to $(V_{s}^{x})_{s\in[0,t]}$, $x\in\mathbb{R}$.
@@
[…]
@@
$K_{s}$ $=-\eta_{s}=-L_{s}+\sigma_{U,L}s+\sum_{0<r\leq s}\frac{\Delta U_{r}\Delta L_{r}}{1+\Delta U_{r}},$
for $s\in[0,t]$. Since the time-reversed Lévy process $(\tilde{U}_{s},\tilde{L}_{s})_{s\in[0,t]}$ is equal in law to $(-U_{s},-L_{s})_{s\in[0,t]}$ we conclude that $(T_{s},\widetilde{\eta}_{s})_{s\in[0,t]}$ and $(W_{s},K_{s})_{s\in[0,t]}$ are equal in law. The claim follows from Theorems 3.5 and 4.1. ∎
@@
Corollary 4.2 is not surprising, since reversibility of the stochastic flow is intrinsically related to duality, as discussed in various articles, see e.g. [46]. Also in our case, it is possible to deduce Theorem 3.5 directly from Theorem 4.1. To see this, let $(U,L)$, $V^{x}_{s}$ and $R^{y}_{s}$ be as in Corollary 4.2. For $u\in[0,t]$ denote by $\varphi_{u,t}:\mathbb{R}\to\mathbb{R}$ the stochastic flow which transports $x=V_{u}(\omega)$ to $V_{t}(\omega)$, i.e. $\varphi_{u,t}(x)=V_{t}(\omega)$ conditional on $V_{u}=x$ (we suppress the superscript $x$ here and prefer to work with $V_{t}|V_{0}=x$ rather than $V_{t}^{x}$ for the moment). By time-homogeneity of $V$, we have for $s\in[0,t]$
@@
showing that $(R_{s}^{y})_{s\in[0,t]}$, $y\in\mathbb{R}$, is dual to $(V_{s}^{x})_{s\in[0,t]}$, $x\in\mathbb{R}$. By the proof of Corollary 4.2, $(T_{s},\widetilde{\eta}_{s})_{s\in[0,t]}$ is equal in law to $(W_{s},K_{s})_{s\in[0,t]}$, showing that the process given in Theorem 3.5 (call it $\overline{R}_{s}^{y}$ for the moment) is also dual to $R$. Hence we have given another proof of the duality of $(\overline{R}_{s}^{y})$ to $(V_{s}^{x})$ in Theorem 3.5. This last proof is in line with the reasoning given in Sigman and Ryan [46, Cor. 3.1 (1)], who consider $0,\infty)-$valued time-homogeneous Markov processes with certain properties.
Remark 4.4.
According to [26, Prop. 4.4] the Siegmund dual GOU process derived in Theorem 3.5 is even strongly pathwise dual, i.e. both $V$ and $R$ can be realised on a suitable probability space such that
$\{V_{t}^{x}\geq y\}=\{R_{t}^{y}\leq x\}\quad\text{almost surely}.$
This follows from the construction method described in [12], but it is not a direct consequence of our proof, as in (3.7) we use a distributional inequality. Still, using time-reversals, one could replace (3.7) by an almost sure statement (see [34, Lemma 6.1] for a special case), and the resulting dual in this case turns out to coincide with the inverse stochastic flow obtained in Theorem 4.1.
Acknowledgements
The authors thank two anonymous referees whose comments and suggestions significantly helped to improve the papers overall presentation and led to additional results such as Corollary 3.10.